OpenStudy (anonymous):

I need to factor this polynomial using GCF: 4x^(5) + 6x^(-1) -3x^(-4) Any idea how to do this?

OpenStudy (anonymous):

well in this case it wont be the number because 4 6 3 dont have a common. i would factor out the x

OpenStudy (anonymous):

but you first might have to put it into the fraction form first because of the - on the exponents.

OpenStudy (anonymous):

6x^(-1) = 6/x 3x^(-4) = ...? Then factor them

OpenStudy (anonymous):

@Chlorophyll wouldnt the whole term go down too?

OpenStudy (anonymous):

I've got x^(-4) (4x^9 + 6x^3 - 3x) so far. so x to the -4th power is the gcf.

OpenStudy (anonymous):

If the whole term goes down, it should be (6x)^ (-1)

OpenStudy (anonymous):

ok i see your point.

OpenStudy (anonymous):

@Chlorophyll that method doesnt make any sense why is it that i need to make an equation when i am using greatest common factor?

OpenStudy (anonymous):

No, I separate them out so it's easy for you to do it one by one Then combine them at the end!

OpenStudy (anonymous):

what if we raised it by the 4th power and simplify from there?

OpenStudy (anonymous):

@hereweg0 isnt simplifying not the same as factoring. Ive made that mistake before.

OpenStudy (anonymous):

well i meant factor.

OpenStudy (anonymous):

or would it not be the same answer if we raised it? @Chlorophyll ? @amelapal? or am i speaking nonsense?

OpenStudy (anonymous):

what you're saying isnt nonsense, its just confusing. what are we raising to the 4th power?

OpenStudy (anonymous):

@hereweg0

OpenStudy (anonymous):

the whole polynomial. but will it work?

OpenStudy (anonymous):

Nope because we are raising it before we even found the greatest common factor.

OpenStudy (anonymous):

right. is it safe to say it is the most factored form?

OpenStudy (anonymous):

I factored it already but i don't know what to do after  is that at its most factored form or do we turn it into a fraction because of the one negative exponent?